create a batch file program that performs 4 mathematical operations such as addition, subtraction, multiplication and division.
1
Expert's answer
2013-02-19T07:56:51-0500
@echo off


set /p a=Enter a:


set /p b=Enter b:


set /a sum=%a%+%b%


echo a + b = %sum%


set /a difference=%a%-%b%


echo a - b = %difference%


set /a product=%a%*%b%


echo a * b = %product%


set /a quotient=%a%/%b%


echo a / b = %quotient%


pause